Sarah and Keith purchased 40 stamps, a mixture of .44 and .20 cent stamps. Find the number of each type of stamp if they spent $16.40.
Answer by checkley79(3341)

X+Y=40
X=40-Y
.44X+.20Y=16.40
.44(40-Y)+.20Y=16.40
17.60-.44Y+.20Y=16.40
-.24Y=16.40-17.60
-.24Y=-1.20
Y=-1.20/-.24
Y=5 $.20 STAMPS.
40-5=45 $.44 STAMPS.
PROOF:
.44*35+.20*5=16.40
15.40+1.00=16.40
16.40=16.40